# SPORTU: A Comprehensive Sports Understanding Benchmark for Multimodal Large Language Models

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) are advancing the ability to reason about complex sports scenarios by integrating textual and visual information. To comprehensively evaluate their capabilities, we introduce SPORTU, a benchmark designed to assess MLLMs across multi-level sports reasoning tasks. SPORTU comprises two key components: SPORTU-text, featuring 900 multiple-choice questions with human-annotated explanations for rule comprehension and strategy understanding. This component focuses on testing models’ ability to reason about sports solely through question-answering (QA), without requiring visual inputs; SPORTU-video, consisting of 1,701 slow-motion video clips across 7 different sports and 12,048 QA pairs, designed to assess multi-level reasoning, from simple sports recognition to complex tasks like foul detection and rule application. We evaluate four prevalent LLMs mainly utilizing few-shot learning paradigms supplemented by chain-of-thought (CoT) prompting on the SPORTU-text part. We evaluate four LLMs using few-shot learning and chain-of-thought (CoT) prompting on SPORTU-text. GPT-4o achieves the highest accuracy of 71%, but still falls short of human-level performance, highlighting room for improvement in rule comprehension and reasoning. The evaluation for the SPORTU-video part includes 7 proprietary and 6 open-source MLLMs. Experiments show that models fall short on hard tasks that require deep reasoning and rule-based understanding. Claude-3.5-Sonnet performs the best with only 52.6% accuracy on the hard task, showing large room for improvement. We hope that SPORTU will serve as a critical step toward evaluating models’ capabilities in sports understanding and reasoning. The dataset is available at https://github.com/haotianxia/SPORTU.
@@ 1 Introduction
To gain deeper insights into SPORTU, we evaluate 4 LLMs on SPORTU-text and 13 MLLMs that handle videos or image sequences on SPORTU-video, including both open-source and proprietary models. We selected GPT-4o, Claude-3.5, LLaMA-3, and Gemini 1.5 Pro for SPORTU-text, while for SPORTU-video, we included a broader range such as GPT-4V, Gemini 1.5 Flash, and Video-ChatGPT. The models were evaluated using few-shot learning for SPORTU-text and multi-frame input for SPORTU-video. Key results show that GPT-4o performs best on SPORTU-text, achieving 71% accuracy.Claude-3.5-Sonnet leads the SPORTU-video evaluation with an overall accuracy of 69.52%, but it achieves only 52.57% on hard-level tasks, indicating challenges with more complex reasoning.
In summary, our contributions are threefold:
  • We introduce SPORTU, a comprehensive Sports Understanding Benchmark that integrates both text-based and video-based tasks to evaluate models’ sports reasoning and knowledge application capabilities. SPORTU-text tests models’ understanding of facts, rules, strategies, and scenario-based sports knowledge, while SPORTU-video presents a three-tiered difficulty level, featuring challenging sports scenarios such as foul detection and complex rule application, with slow-motion footage to capture fine-grained details essential for understanding nuanced actions and rule violations.
  • This research deepens our understanding of LLM and MLLM performance on sports tasks, showing GPT-4o achieved 71% accuracy on SPORTU-text, while Claude-3.5-Sonnet led SPORTU-video with 69.52% overall accuracy. However, Claude-3.5-Sonnet struggled on hard-level tasks, achieving only 52.57%, highlighting significant gaps in deep reasoning with rule comprehension.
  • We will make the SPORTU benchmark publicly available to support advancements in sports analytics and applications.
@@ 2.1 Multimodal Sports Analysis
Table 1: Comparison of various QA datasets across multiple dimensions. This table highlights key distinctions between SPORTU and other QA datasets, including QA Type (MC: Multiple Choice, OE: Open Ended), whether explanations are provided, the presence of Multilevel Difficulty, and coverage of Domain Knowledge in terms of both Action and Rule. Additional features compared include whether datasets incorporate Slow Motion, Multi-Camera Angles, and the number of sports covered. Avg Exp. or Ans. Length refers to the average word count for open-ended explanations or answers.
Benchmark QA Type Question Type Explanation Multilevel Diff. Domain Knowledge Slow Motion Multi Camera Angle No. of Sports Avg Exp. or Ans. Length Action Rule Text BIG-bench Context-free MC - - 5 - LiveQA Context extractive MC - - 1 - QASports Context extractive OE - - 3 - SportQA Context-free MC - - 35 - SPORTU-text (Ours) Context-free MC - - 5 100.19 Video Sports-QA Video QA OE 8 - SoccerNet-XFoul Video QA OE 1 25 SPORTU-video (Ours) Video QA OE + MC 7 13.29
@@ 2.2 Multimodal Sports QA
Question answering (QA) has been widely adopted to evaluate the comprehension abilities of LLMs across various domains. However, existing QA datasets have limited coverage of sports understanding, as shown in 1. Several QA datasets have been introduced to evaluate models’ understanding of textual information, ranging from factual recall to multi-hop reasoning (Joshi et al., 2017; Yang et al., 2018; Clark et al., 2019). In the multimodal domain, QA Benchmarks have been extended to evaluate MLLMs’ image and video understanding by answering questions based on visual inputs (Fu et al., 2024; Yue et al., 2024a; b; He et al., 2024a; Zhou et al., 2024). However, sports-related QA is underrepresented in these general datasets, and even when sports topics are included, the questions often lack sufficient difficulty due to the datasets’ broader focus. Existing general QA datasets tend to focus on surface-level aspects of sports, such as historical facts or well-known events in text-based tasks, and perception-level tasks, like object or action identification in video-based tasks. As a result, they often fail to assess the deep, domain-specific knowledge and reasoning required for a nuanced understanding of sports.
@@
On the other hand, sports-domain-specific Video QA(VQA) datasets that effectively combine video and text modalities to test comprehensive sports understanding are scarce. Sports-QA (Li et al., 2024a) stands out by covering eight sports, including volleyball and basketball, with videos sourced from MultiSports (Li et al., 2021) and FineGym (Shao et al., 2020). However, although it provides detailed action recognition annotations, Sports-QA does not assess models on understanding sports rules, such as foul detection. This limitation arises because MultiSports does not include foul clips, and FineGym focuses solely on granular gymnastic actions, which do not encompass rule-based scenarios.
@@ 3 Sports Understanding Benchmark
SPORTU-video – the first slow-motion VQA dataset covering multiple sports, designed to evaluate MLLMs’ sports understanding abilities. It features a multilevel question design using slow-motion videos, with some clips offering multiple camera angles. The tasks range from easy-level tasks, such as sports recognition, to medium-level tasks, such as team role recognition, and hard-level tasks, such as rule violation explanations. The dataset contains 1,701 slow-motion video clips across 7 sports, with 12,048 QA pairs designed to test models’ multi-level video-text reasoning capabilities.
@@
To guarantee the accuracy of annotations in SPORTU, we employed a rigorous quality verification process. Our annotator team consisted of nine sports experts, including two intercollegiate athletes and seven players with over five years of training experience (the detailed annotation and verification process is described in Appendix A).
@@ 3.1 SPORTU-text: Pure Text QA
Dataset Construction Among all sports-specific QA benchmarks, only SportQA (Xia et al., 2024a) Level-3 questions assess models’ deep understanding of sports knowledge. We randomly selected 900 questions from five different sports: American football, basketball, soccer, volleyball, and tennis. Each question was carefully annotated, with detailed explanations provided for each option—whether correct or incorrect. Annotators explained why an option was correct or not, offering clear insights into the reasoning required for each answer. Examples can be found in the appendix E or in the actual SPORTU-text dataset.
@@ 3.2 SPORTU-video: Multimodal Video QA
SPORTU-video consists of 1,701 slow-motion video clips across 7 different sports, including soccer, basketball, volleyball, ice hockey, tennis, baseball, and badminton. We also ensured that for some sports, the videos featured multiple camera angles to challenge the models’ ability to capture consistent judgments across different perspectives. Specifically, 300 video scenes were selected to include multi-angle views. Each clip is accompanied by one or more QA pairs, for a total of 12,048 QA pairs (with 10,973 Multiple Choice Questions and 1,075 open-ended questions based on the explanations that the annotators labeled), with three levels of complexity: Easy: 25.36%, Medium: 50.22%, Hard: 24.42%.
Dataset Construction The construction of SPORTU-video began by identifying the types of tasks that could be asked based on the sports domain. Some questions, such as sports recognition and identifying the number of players, are common across all sports. Other questions are sports-specific, requiring knowledge of rules or strategies unique to each sport. The full question template can be found in the appendix D.
@@
To classify the questions, we considered three different levels of difficulty based on the sports knowledge and reasoning required. Easy-level questions: These tasks rely on commonsense, such as basic sports recognition. For example, questions like ”What sport does the video show?” or ”How many players are shown?” do not require sports knowledge. Medium-level questions: These questions require sports knowledge beyond commonsense. For example, models are asked to identify which team is on offense based on the video or to recognize specific roles, such as a libero in volleyball, by identifying the libero’s jersey color. Hard-level questions: These tasks involve deep rule-based reasoning. For example, identifying rule violations or technical errors requires models to understand the sport’s rules in detail and apply them to the specific context of the video.
@@
Once the question types were defined, the annotators collected the appropriate video clips and labeled the corresponding ground truths that the videos showed, as multiple questions could often be answered from a single video. These slow-motion clips were sourced from YouTube under the Creative Commons License, and each video was manually cropped to ensure high-quality footage suitable for detailed action analysis and rule comprehension. For each video, annotators carefully reviewed the content and annotated the answers for the designed questions. After the initial annotation, we used templates to generate both open-ended questions and multiple-choice questions based on the annotator-labeled ground truths. For multiple-choice questions, the ground truth answer was included as one of the options, while distractors were selected based on the question type. For example, distractors for jersey color questions were drawn from a set of plausible colors, while for ”what sport” questions, the distractors consisted of similar sports. In foul-related questions, the distractors included other types of fouls within the same sport. For open-ended questions, the annotators’ explanations were used directly as the ground truth.
@@
4 Experiments
We compare MLLM’s performance on the SPORTU benchmark, We also evaluate the ability of models to produce explanations. We start by describing the MLLMs in our experiments and their Experimental Setup(§4.1), followed by evaluation metrics (§4.2).
4.1 Experimental Setup
@@
We primarily focused on the Chain-of-Thought (CoT) prompting method, based on evidence from Wei et al. (2022), which demonstrates the effectiveness of few-shot CoT in reasoning tasks, including sports understanding. In addition to CoT, we explored standard prompting (SP) for both zero-shot and few-shot settings (Brown, 2020) and zero-shot CoT (Kojima et al., 2022) as baselines. In the few-shot settings, five exemplars were provided for each task, with answers annotated by human experts. Exemplars were consistently sourced from the development set, with the temperature parameter set to 0 to ensure consistent response generation.
SPORTU-video Evaluation: We investigate a range of MLLMs, including 7 close source models and 6 open source models. For close source models, we evaluate GPT-4o (2024-08-06 version) and -4omini (OpenAI, 2024a), GPT-4V (OpenAI, 2023), Gemini 1.5 Pro (Gemini Team, 2024a), Gemini 1.5 Flash (Gemini Team, 2024b) and Claude-3.5-Sonnet (20240620 version) (Anthropic, 2024b) and-3.0-Haiku Anthropic (2024).For Open source models, we evaluate ChatUniVi (Jin et al., 2023), LLaVA-NeXT (Liu et al., 2024), mPLUG-Owl3 (Ye et al., 2024), Tarsier (Wang et al., 2024), Video-ChatGPT (Maaz et al., 2024), and VideoChat2(Li et al., 2024b). For the closed-source models, we adhered to the default settings provided by their official APIs. GPT and Claude family models processed ten image frames extracted from the video content as input. The Gemini family models processed the entire video, as their API supports video input. Due to computing resource limitations, we used the 7b versions of all open-source models in this evaluation. For VideoChat, we set ‘max_frames’ to 100, while for the other open-source models, we used 16 frames as input. Across all closed and open-source models, we set the temperature parameter to 0 to ensure consistent response generation.
Models were prompted to generate the answer first, followed by an explanation for why the model chose that answer. This method has been shown to be effective in generating interpretable results in reasoning tasks (Chen et al., 2023), and it also facilitates human error analysis. More details regarding the specific prompts used for both SPORTU-text and SPORTU-video can be found in Appendix B. All inferences are run on an NVIDIA RTX A6000.
@@ 4.2 Evaluation Metrics
We use accuracy (prediction compared to ground truth) to evaluate the predictions of each model in multiple-choice tasks. We explore several methods to evaluate different aspects of model-generated explanations and open-ended questions: ROUGE-L (Lin, 2004), BERTScore (Zhang et al., 2019), BLEURT (Sellam et al., 2020), CTC-Preservation (Deng et al., 2021), and G-Eval Score (Liu et al., 2023). ROUGE-L computes the surface-form similarity between model-generated explanations and reference (human-annotated) explanations. BERTScore and BLEURT measure semantic similarity using pre-trained BERT (Devlin, 2018) and fine-tuned BERT models respectively. CTC metrics evaluate the information alignment of model-generated explanations. G-Eval is a framework that uses large language models with chain-of-thought reasoning and a form-filling paradigm to assess the quality of NLG outputs. We apply this framework to evaluate the accuracy, conciseness, and relevance of the model-generated explanations compared to the ground truth, reporting these as an overall score. We use GPT-4o to implement the G-Eval process. The detailed prompt can be found in the Appendix B. To ensure the reliability of the G-Eval scores, we conducted a human evaluation on a randomly selected set of 20 questions per sport across 7 sports, totaling 140 questions. This subset was used to compare the G-Eval scores with human-annotated scores across 13 models. By using the same set of questions for all models, we ensured consistency in comparison, allowing us to assess whether the G-Eval scores were in line with human judgment. The human annotators rated the model-generated content using the same criteria as G-Eval, providing a reference point for how well the automated evaluation aligns with human evaluation.

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) are advancing the ability to reason about complex sports scenarios by integrating textual and visual information. To comprehensively evaluate their capabilities, we introduce SPORTU, a benchmark designed to assess MLLMs across multi-level sports reasoning tasks. SPORTU comprises two key components: SPORTU-text, featuring 900 multiple-choice questions with human-annotated explanations for rule comprehension and strategy understanding. This component focuses on testing models’ ability to reason about sports solely through question-answering (QA), without requiring visual inputs; SPORTU-video, consisting of 1,701 slow-motion video clips across 7 different sports and 12,048 QA pairs, designed to assess multi-level reasoning, from simple sports recognition to complex tasks like foul detection and rule application. We evaluated four prevalent LLMs mainly utilizing few-shot learning paradigms supplemented by chain-of-thought (CoT) prompting on the SPORTU-text part. GPT-4o achieves the highest accuracy of 71%, but still falls short of human-level performance, highlighting room for improvement in rule comprehension and reasoning. The evaluation for the SPORTU-video part includes 6 proprietary and 8 open-source MLLMs. Experiments show that models fall short on hard tasks that require deep reasoning and rule-based understanding. GPT-4o performs the best with only 57.8% accuracy on the hard task, showing large room for improvement. We hope that SPORTU will serve as a critical step toward evaluating models’ capabilities in sports understanding and reasoning. The dataset is available at https://github.com/haotianxia/SPORTU.
@@ 1 Introduction
To gain deeper insights into SPORTU, we evaluated 4 LLMs on SPORTU-text and 14 MLLMs on SPORTU-video, covering both open-source and proprietary models. For SPORTU-text, we selected GPT-4o, Claude-3.5-Sonnet, LLaMA-3.1, and Gemini 1.5 Pro, using few-shot learning paradigms. For SPORTU-video, we evaluated a broader range of models, including GPT-4o, Gemini 1.5 Flash, and Video-ChatGPT, using multi-frame inputs.
Key results reveal that GPT-4o achieves the highest accuracy on SPORTU-text at 71%, highlighting its relatively strong performance in text-based sports reasoning. On SPORTU-video, Qwen2-VL 72B achieves the highest overall accuracy of 70.94%, but struggles on hard-level tasks with only 44.12% accuracy. Claude-3.5-Sonnet demonstrates the best performance on hard-level tasks (52.57%), yet the results across all models reveal significant challenges in handling complex reasoning and rule comprehension, particularly in tasks requiring deep sports knowledge. We also systematically applied different prompt strategies to evaluate reasoning performance in SPORTU-video. The results reveal that models generally achieve the highest overall performance when directly predicting the answer without providing a rationale. However, when models are required to generate a rationale first and then predict the answer, performance declines. For instance, Claude-3.5-Sonnet’s accuracy dropped from 52.57% with direct answer prediction to 39.32% when reasoning the rationale first. This indicates that current models struggle to maintain consistency and robustness in reasoning for complex tasks. These findings highlight the need for future advancements in reasoning capabilities.
@@ 2.2 Multimodal Sports QA
Question answering (QA) has been widely adopted to evaluate the comprehension abilities of LLMs across various domains. Several QA datasets have been introduced to evaluate models’ understanding of textual information, ranging from factual recall to multi-hop reasoning (Joshi et al., 2017; Yang et al., 2018; Clark et al., 2019). In the multimodal domain, QA Benchmarks have been extended to evaluate MLLMs’ image and video understanding by answering questions based on visual inputs (Fu et al., 2024; Yue et al., 2024a; b; He et al., 2024a; Zhou et al., 2024). However, sports-related QA is underrepresented in these general datasets, and even when sports topics are included, the questions often lack sufficient difficulty due to the datasets’ broader focus. Existing general QA datasets tend to focus on surface-level aspects of sports, such as historical facts or well-known events in text-based tasks, and perception-level tasks, like object or action identification in video-based tasks. As a result, they often fail to assess the deep, domain-specific knowledge and reasoning required for a nuanced understanding of sports.
@@
On the other hand, sports-domain-specific VQA datasets that effectively combine video and text modalities to test comprehensive sports understanding are scarce. Sports-QA (Li et al., 2024a) stands out by covering eight sports, including volleyball and basketball, with videos sourced from MultiSports (Li et al., 2021) and FineGym (Shao et al., 2020). However, although it provides detailed action recognition annotations, Sports-QA does not assess models on understanding sports rules, such as foul detection. This limitation arises because MultiSports does not include foul clips, and FineGym focuses solely on granular gymnastic actions, which do not encompass rule-based scenarios.
@@ 3 Sports Understanding Benchmark
SPORTU-video – the first VQA dataset covering multiple sports, designed to evaluate MLLMs’ sports understanding abilities. It features a multilevel question design using slow-motion videos, with some clips offering multiple camera angles. The tasks range from easy-level tasks, such as sports recognition, to medium-level tasks, such as Team Role Recognition, and hard-level tasks, such as rule violation explanations. The dataset contains 1,701 slow-motion video clips across 7 sports, with 12,048 QA pairs designed to test models’ multi-level video-text reasoning capabilities.
@@
3.1 Quality Control
@@ 3.2 SPORTU-text: Pure Text QA
Dataset Construction Among all sports-specific QA benchmarks, only SportQA (Xia et al., 2024a) Level-3 questions assess models’ deep understanding of sports knowledge by covering rule-, strategy-, and scenario-based questions. However, these questions are mixed together without clear labels for each type. To build SPORTU-text, we randomly selected 900 questions from five different sports: American football, soccer, volleyball, basketball, and tennis, and our expert annotators manually categorized the selected questions. Rule-related questions focus on explicit sports rules, strategy-related questions involve tactical or strategic decisions, and scenario-related questions provide a specific context or player interaction (e.g., “Player A performs action X, and Player B reacts”). While some questions could fit multiple categories, annotators assigned each question to the category that most accurately reflected its primary focus. Each question was carefully annotated, with detailed explanations provided for each option—whether correct or incorrect. Annotators explained why an option was correct or not, offering clear insights into the reasoning required for each answer. This additional annotation step ensures a more structured dataset and enables the evaluation of models’ reasoning capabilities across distinct aspects of sports knowledge. Examples can be found in the appendix O.
@@ 3.3 SPORTU-video: Multimodal Video QA
SPORTU-video consists of 1,701 slow-motion video clips across 7 different sports, including soccer, basketball, volleyball, ice hockey, tennis, baseball, and badminton. We also ensured that for some sports, the videos featured multiple camera angles to challenge the models’ ability to capture consistent judgments across different perspectives. Our expert annotators manually cropped video clips from replay footage to include multiple perspectives of the same foul, as such replays are standard in sports broadcasts. This process ensures accuracy while minimizing additional manual work. Specifically, 300 video scenes were selected to include multi-angle views. Each clip is accompanied by one or more QA pairs, for a total of 12,048 QA pairs (with 10,973 Multiple Choice Questions and 1,075 open-ended questions based on the explanations that the annotators labeled), with three levels of complexity: Easy: 25.36%, Medium: 50.22%, Hard: 24.42%.
@@
Dataset Construction The construction of SPORTU-video began by identifying the types of tasks that could be asked based on the sports domain. Some questions, such as sports recognition and identifying the number of players, are common across all sports. Other questions are sports-specific, requiring knowledge of rules or strategies unique to each sport. Full question templates can be found in the appendix N. To classify the questions, we considered three different levels of difficulty based on the sports knowledge and reasoning required. Easy-level questions: These tasks rely on commonsense, such as basic sports recognition. For example, questions like “What sport does the video show?” or “How many players are shown?” do not require sports knowledge. Medium-level questions: These questions require sports knowledge beyond commonsense. For example, models are asked to identify which team is on offense based on the video or to recognize specific roles, such as a libero in volleyball, by identifying the libero’s jersey color. Hard-level questions: These tasks involve deep rule-based reasoning. For example, identifying rule violations or technical errors requires models to understand the sport’s rules in detail and apply them to the specific context of the video.
@@
Once the question types were defined, the annotators collected the appropriate video clips and labeled the corresponding ground truths that the videos showed, as multiple questions could often be answered from a single video. These slow-motion clips were sourced from YouTube under the Creative Commons License, and each video was manually cropped to ensure high-quality footage suitable for detailed action analysis and rule comprehension. For multiple-choice questions, annotators labeled the ground truth category, such as the specific foul (“handball” or “offsid”), which was then used to generate multiple-choice questions with distractors derived from other categories. For open-ended questions, annotators provided detailed explanations for the rule violation or foul observed in the video. These explanations were used as the ground truth for generating open-ended questions, allowing models to be tested on reasoning and explanation generation.
@@
4 Experiment
We compare MLLM’s performance on the SPORTU benchmark, We also evaluate the ability of models to produce explanations. We start by describing the MLLMs in our experiments and their experimental settings(§4.1), followed by prompting strategies (§4.2), and evaluation metrics (§4.3).
4.1 Models
@@
SPORTU-video Evaluation: We investigate a range of MLLMs, including 6 close-source models and 6 open-source models. For close-source models, we evaluated GPT-4o (2024-08-06 version) and -4omini (OpenAI, 2024a), Gemini 1.5 Pro (Gemini Team, 2024a), Gemini 1.5 Flash (Gemini Team, 2024b), Claude-3.5-Sonnet (20240620 version) (Anthropic, 2024b) and-3.0-Haiku Anthropic (2024).For open-source models, we evaluated ChatUniVi (Jin et al., 2023), LLaVA-NeXT (Liu et al., 2024), mPLUG-Owl3 (Ye et al., 2024), Tarsier (Wang et al., 2024), Video-ChatGPT (Maaz et al., 2024), VideoChat2(Li et al., 2024b), ST-LLM (Liu et al., 2025), and Qwen2-VL-72B (Bai et al., 2023). For the closed-source models, we adhered to the default settings provided by their official APIs. GPT and Claude family models processed ten image frames extracted from the video content as input. The Gemini family models processed the entire video, as their API supports video input. Due to computing resource limitations, we used the 7B versions of all open-source models in this evaluation except Qwen2-VL-72B, which can be accessed by API. For VideoChat, we set ‘max_frames’ to 100, while for the other open-source models, we used 16 frames as input. Across all closed and open-source models, we set the temperature parameter to 0 to ensure consistent response generation. All inferences are run on an NVIDIA RTX A6000.
4.2 Prompting Strategies
We apply three different prompting strategies to generate answers and/or explanations. We represent the input as X for the question and answer options, Y for the answer, and R for the explanation (rationale). Our three strategies are:
  • $X\rightarrow Y$: No-CoT, which directly predicts the answer.
  • $X\rightarrow RY$: Reasoning where answer inference is conditioned to the rationale. This strategy asks the model to engage in step-by-step reasoning first, and then answer the question. This approach is based on chain of thought (CoT) prompting, which has been shown to improve LLMs’ prediction accuracy across various reasoning tasks Wei et al. (2022). The zero-shot CoT method is adapted from Kojima et al. (2022). We prompt the model to ‘think step by step, then provide the correct answer.’
  • $X\rightarrow YR$: This strategy asks the model to answer the question first, followed by the rationale for why the model chose that option. This prompting method has proven effective on REV Chen et al. (2023).
For the SPORTU-text evaluation, LLMs have demonstrated their capability for in-context learning by utilizing exemplars through few-shot prompting (Brown, 2020). Therefore, we use four prompting baselines for evaluation: zero-shot $X\rightarrow Y$ (0S), zero-shot $X\rightarrow RY$ (0CoT), five-shot $X\rightarrow Y$ (5S), and five-shot $X\rightarrow RY$ (5CoT). For the five-shot method, we provide five exemplars with only the answers. The five exemplars for the five-shot CoT method include both the answers and human-annotated rationales.
For the SPORTU-video evaluation, we use all three prompting methods. The zero-shot $X\rightarrow YR$ prompting method is applied to conduct human error analysis. More details of models’ specific prompts are shown in the appendix L.
@@ 4.3 Evaluation Metrics
We use accuracy (prediction compared to ground truth) to evaluate the predictions of each model in multiple-choice tasks. We explore several methods to evaluate different aspects of model-generated explanations and open-ended questions: ROUGE-L (Lin, 2004), BERTScore (Zhang et al., 2019), BLEURT (Sellam et al., 2020), CTC-Preservation (Deng et al., 2021), and G-Eval Score (Liu et al., 2023). ROUGE-L computes the surface-form similarity between model-generated explanations and reference (human-annotated) explanations. BERTScore and BLEURT measure semantic similarity using pre-trained BERT (Devlin, 2018) and fine-tuned BERT models respectively. CTC metrics evaluate the information alignment of model-generated explanations. G-Eval is a framework that uses large language models with chain-of-thought reasoning and a form-filling paradigm to assess the quality of NLG outputs. We apply this framework to evaluate the accuracy, conciseness, and relevance of the model-generated explanations compared to the ground truth, reporting these as an overall score. Since LLMs might favor their own answers, we utilize all four models evaluated in the SPORTU-text part to implement the G-Eval process for SPORTU-text results and three models (excluding LLama3.1-405B) for the SPORTU-video part. We also calculate the average score of the models to obtain an overall score. The detailed prompt can be found in the Appendix J. To ensure the reliability of the G-Eval scores, we conducted a human evaluation on a randomly selected set of 20 questions per sport across 7 sports, totaling 140 questions. This subset was used to compare the G-Eval scores with human-annotated scores across 14 models. By using the same set of questions for all models, we ensured consistency in comparison, allowing us to assess whether the G-Eval scores were in line with human judgment. The human annotators rated the model-generated content using the same criteria as G-Eval, providing a reference point for how well the automated evaluation aligns with human evaluation.